The equation for the regression line modeling the relationship between test scores and sleep time is...

  1. The equation for the regression line modeling the relationship between test scores and sleep time is ???̂? ????? = 63 + 3.4h???? ?????, with r=0.77. If a student sleeps just 3 hours before the exam, what score are they expected to get? Write your answer in a complete sentence.
Answer #1

Our regression equation is:

Test score = 63 + 3.4*hours of sleep

So the slope coefficient is 3.4 (ie when hours of sleep increases by 1, test score increases by 3.4)

Intercept coefficient is 63, ie when hours of sleep is 0, test score will still be 63.

If we put, hours of sleep = 3 in the above equation, we get, expected test score = 63 + 3.4*3 = 73.2